# Geometric calculations on probability manifolds from reciprocal relations in Master equations

### v4

@@ 3.2. Riemannian curvature tensor
$\begin{split}&\frac{1}{2}\sum_{(i,j)\in E}(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}(\nabla_{p}\theta L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{ij}\\ =&\frac{1}{4}\sum_{(i,j)\in E}(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\Big(\frac{\partial\theta_{ij}}{\partial p_{i}}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{i}+\frac{\partial\theta_{ij}}{\partial p_{j}}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{j}\Big)\\ =&\frac{1}{2}\sum_{(i,j)\in E}(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{i}\\ =&\frac{1}{2}\sum_{i=1}^{n}\Gamma(\Phi_{2},\Phi_{4},p)_{i}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{i}\\ =&\frac{1}{4}\sum_{i,j=1}^{n}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}(\mathsf{V}_{\Phi_{1}}\theta)_{ij}\\ =&\frac{1}{2}\sum_{i,j=1}^{n}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}(L(\theta)\Phi_{1})_{i}\\ =&\frac{1}{2}\sum_{i,j=1}^{n}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}\sum_{k=1}^{n}\sqrt{\omega_{ik}}(\nabla_{\omega}\Phi_{1})_{ki}\theta_{ik}\\ =&\frac{1}{4}\sum_{i,j,k=1}^{n}(\nabla_{\omega})_{ik}\Big((\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}\Big)(\nabla_{\omega}\Phi_{1})_{ik}\theta_{ik}\\ =&\sum_{(i,j)\in E}\frac{1}{2}\Gamma^{3}(\Phi_{2},\Phi_{4},\Phi_{1},\Phi_{3},p)_{ij}\theta_{ij}.\end{split}$
Thirdly, we switch the indices $1$ and $3$ in the above formula. We last derive
@@ 4.2. Levi-Civita connection in probability manifolds
$\begin{split}\Gamma(\Phi_{1},\Phi_{2},p)=&\Big((\nabla_{\omega}\Phi_{1})_{12}(\nabla_{\omega}\Phi_{2})_{12}\frac{\partial\theta_{12}}{\partial p_{i}}+(\nabla_{\omega}\Phi_{1})_{13}(\nabla_{\omega}\Phi_{2})_{13}\frac{\partial\theta_{13}}{\partial p_{i}}\\ &+(\nabla_{\omega}\Phi_{1})_{23}(\nabla_{\omega}\Phi_{2})_{23}\frac{\partial\theta_{23}}{\partial p_{i}}\Big)_{i=1}^{3}.\end{split}$
Similarly, one can define $\Gamma(\Phi_{1},\Phi_{3},p),\Gamma(\Phi_{2},\Phi_{3},p)\in\mathbb{R}^{3}$. Then the commutator in $(\Delta_{3},\mathbf{g})$ formulates
@@ 5. Example II: Curvatures of probability manifolds on a three-point lattice graph
where $\mathsf{V}_{\Phi}=L(\theta)\Phi$, and we let $\omega_{12}=\omega_{23}=1$, $\omega_{13}=0$, and
@@
There is a particular coordinate for $\Delta_{3}$, which simplifies the geometric calculations. Denote the cumulative distribution function (CDF) on discrete states, such that
@@
Denote a set $\mathrm{CDF}=\{(x_{1},x_{2})\in[0,1]^{2}\colon x_{1}\leq x_{2}\}$. Thus, the metric $\mathbf{g}$ in coordinates $(x_{1},x_{2})$ is a diagonal metric, such that $\mathbf{g}=(\mathbf{g}_{ij})_{1\leq i,j\leq 2}\in\mathbb{R}^{2\times 2}$, with
@@
$\bar{R}_{11}(p)=\bar{K}_{12}(p)\theta_{2}(p),\quad\bar{R}_{22}(p)=\bar{K}_{12}(p)\theta_{1}(p),\quad\bar{R}_{12}(p)=0.$
@@
The proof is based on standard calculations of a diagonal metric. The Riemannian curvature tensor $\bar{R}$ in $(\mathrm{CDF},\mathbf{g})$ is defined below. Denote $i,j,k,l=1,2$, and write $\partial_{i}=\frac{\partial}{\partial x_{i}}$. Then,
$\bar{R}^{i}_{jkl}(x)=\bar{R}(\frac{\partial}{\partial x_{k}},\frac{\partial}{\partial x_{l}})\frac{\partial}{\partial x_{i}}.$
Denote the Levi-Civita connection by
@@
From the definition of curvatures, we have that the section curvature is defined by $\bar{K}_{12}=\frac{1}{\theta_{2}}\bar{R}_{121}^{2}$, the Ricci tensor is defined by $\bar{R}_{jk}=\sum_{i=1}^{2}\bar{R}^{i}_{kij}$, and the scalar curvature is defined by $\bar{S}=\sum_{i=1}^{2}\theta_{i}\bar{R}_{ii}$. We then exchange the coordinates from the cumulative distribution function $x$ to the probability function $p$. This finishes the proof. ∎
@@
We present the proofs of Example 4 and 5 in the appendix.
@@ 6. Discussions
Acknowledgements. W. Li’s work was supported by the AFOSR YIP Award (FA9550-23-1-0087), NSF RTG (2038080), NSF DMS (2245097), and the McCausland Faculty Fellowship at the University of South Carolina.

### v5

@@ 3.2. Riemannian curvature tensor
$\begin{split}&\frac{1}{2}\sum_{(i,j)\in E}(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}(\nabla_{p}\theta L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{ij}\\ =&\frac{1}{4}\sum_{(i,j)\in E}(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\Big(\frac{\partial\theta_{ij}}{\partial p_{i}}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{i}+\frac{\partial\theta_{ij}}{\partial p_{j}}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{j}\Big)\\ =&\frac{1}{2}\sum_{(i,j)\in E}(\nabla_{\omega}\Phi_{2})_{ij}(\nabla_{\omega}\Phi_{4})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{i}\\ =&\frac{1}{2}\sum_{i=1}^{n}\Gamma(\Phi_{2},\Phi_{4},p)_{i}(L(\mathsf{V}_{\Phi_{1}}\theta)\Phi_{3})_{i}\\ =&\frac{1}{4}\sum_{i,j=1}^{n}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}(\mathsf{V}_{\Phi_{1}}\theta)_{ij}\\ =&\frac{1}{2}\sum_{i,j=1}^{n}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}(L(\theta)\Phi_{1})_{i}\\ =&\frac{1}{2}\sum_{i,j=1}^{n}(\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}\sum_{k=1}^{n}\sqrt{\omega_{ik}}(\nabla_{\omega}\Phi_{1})_{ik}\theta_{ik}\\ =&\frac{1}{4}\sum_{i,j,k=1}^{n}(\nabla_{\omega})_{ik}\Big((\nabla_{\omega}\Gamma(\Phi_{2},\Phi_{4},p))_{ij}(\nabla_{\omega}\Phi_{3})_{ij}\frac{\partial\theta_{ij}}{\partial p_{i}}\Big)(\nabla_{\omega}\Phi_{1})_{ik}\theta_{ik}\\ =&\sum_{(i,j)\in E}\frac{1}{2}\Gamma^{3}(\Phi_{2},\Phi_{4},\Phi_{1},\Phi_{3},p)_{ij},\end{split}$
where we apply the definition of $\Gamma^{3}$. Thirdly, we switch the indices $1$ and $3$ in the above formula. We last derive
@@ 4.2. Levi-Civita connection in probability manifolds
$\begin{split}\Gamma(\Phi_{1},\Phi_{2},p)=&\Big((\nabla_{\omega}\Phi_{1})_{12}^{2}\frac{\partial\theta_{12}}{\partial p_{i}}+(\nabla_{\omega}\Phi_{1})_{13}^{2}\frac{\partial\theta_{13}}{\partial p_{i}}+(\nabla_{\omega}\Phi_{1})_{23}^{2}\frac{\partial\theta_{23}}{\partial p_{i}}\Big)_{i=1}^{3}.\end{split}$
Similarly, one can define $\Gamma(\Phi_{1},\Phi_{3},p),\Gamma(\Phi_{2},\Phi_{3},p)\in\mathbb{R}^{3}$. Then the commutator in $(\Delta_{3},\mathbf{g})$ satisfies
@@ 5. Example II: Curvatures of probability manifolds on a three-point lattice graph
where $\mathsf{V}_{\Phi}=L(\theta)\Phi$. We let $\omega_{12}=\omega_{23}=1$, $\omega_{13}=0$, such that
@@
There is a particular coordinate for $\Delta_{3}$, which simplifies geometric calculations. Denote the cumulative distribution function (CDF) on discrete states, such that
@@
Denote a set $\mathrm{CDF}=\{(x_{1},x_{2})\in[0,1]^{2}\colon x_{1}\leq x_{2}\}$. Thus, the metric $\mathbf{g}$ in coordinates $(x_{1},x_{2})$ is a diagonal matrix, such that $\mathbf{g}=(\mathbf{g}_{ij})_{1\leq i,j\leq 2}\in\mathbb{R}^{2\times 2}$, with
@@
$\bar{R}_{11}(p)=\bar{K}_{12}(p)\theta_{2}(p),\quad\bar{R}_{22}(p)=\bar{K}_{12}(p)\theta_{1}(p),\quad\bar{R}_{12}(p)=\bar{R}_{21}(p)=0.$
@@
The proof is based on standard geometric calculations for curvatures with a diagonal metric. The Riemannian curvature tensor $\bar{R}$ in $(\mathrm{CDF},\mathbf{g})$ is computed as follows. Let $i,j,k,l\in\{1,2\}$. Denote the Levi-Civita connection as
@@
We note the following facts from the definition of curvatures. The section curvature is defined by $\bar{K}_{12}=\frac{1}{\theta_{2}}\bar{R}_{121}^{2}$. The Ricci tensor is defined by $\bar{R}_{jk}=\sum_{i=1}^{2}\bar{R}^{i}_{kij}$. And the scalar curvature is defined by $\bar{S}=\sum_{i=1}^{2}\theta_{i}\bar{R}_{ii}$. We then exchange the coordinates from the cumulative distribution function $x$ to the probability function $p$. This finishes the derivation. ∎
@@
In the appendix, we present the proofs of Examples 4 and 5.
@@ 6. Discussions
Acknowledgements. W. Li’s work is supported by AFOSR YIP award No. FA9550-23-1-0087, NSF RTG: 2038080, NSF DMS-2245097, and the McCausland Faculty Fellow in University of South Carolina.